OpenStudy (anonymous):

@ganeshie8 I have a few questions about this problem if you don't mind...? When i'm factoring... say a problem looks like this \[\frac{ 4 }{ m+9 } + \frac{ 5 }{ m^2 - 81}\] \[\frac{ 4 }{ m+9 } + \frac{ 5 }{ (m+9)(m-9) }\] Does the m+9 on the LHS stay the same or can that somehow be factored and do I have to factor the 4 and the 5 or do those stay the same? Or do I get a common denominator and then touch the numerators?

ganeshie8 (ganeshie8):

Since u have two fractions, to add them, they must have same denominators.

ganeshie8 (ganeshie8):

So, multiply and divide first fraction wid \((m-9)\)

ganeshie8 (ganeshie8):

\(\large \frac{ 4 }{ m+9 } + \frac{ 5 }{ (m+9)(m-9) }\) \(\large \frac{ 4 }{ m+9 }\color{green}{\frac{m-9}{m-9}} + \frac{ 5 }{ (m+9)(m-9) }\)

ganeshie8 (ganeshie8):

now both fractions have same denominator, so u can add the numerators

ganeshie8 (ganeshie8):

\(\large \frac{ 4 }{ m+9 } + \frac{ 5 }{ (m+9)(m-9) }\) \(\large \frac{ 4 }{ m+9 }\color{green}{\frac{m-9}{m-9}} + \frac{ 5 }{ (m+9)(m-9) }\) \(\large \frac{ 4(m-9) + 5}{ (m+9)(m-9) }\)

ganeshie8 (ganeshie8):

simplify the numerator

ganeshie8 (ganeshie8):

let me knw if something doesnt make sense :)

OpenStudy (anonymous):

o.o I think it makes sense so far

ganeshie8 (ganeshie8):

good, simplify the numerator and we're done!

OpenStudy (anonymous):

Do I use distributive property and then add 5 to it?

ganeshie8 (ganeshie8):

exactly !

OpenStudy (anonymous):

Um... \[\frac{ 4m - 36 +5 }{ (m+9)(m-9) }\] I'm not sure whether or not I actually add the 5 or leave it alone but if I do add it \[\frac{ 4m - 31 }{ (m-9)(m+9) }\] And thank you for the positive testimonial c:

OpenStudy (amistre64):

are you trying to factor out a m+9, or trying to combine the rational expressions?

ganeshie8 (ganeshie8):

Looks good, and yes always combine like terms and simplify...

OpenStudy (anonymous):

Thank you and I think I'm supposed to be combining rational expressions. I'm not sure... what's the difference between the two if you don't mind me asking?

ganeshie8 (ganeshie8):

np.. u wlc :) there is a slight difference... if you're asked to factor out oly (m+9), then u dont need to simplify further... but here they asked you to combine the rational expressions... so u must factor, get same denominators, add numerators and simplify.

OpenStudy (anonymous):

Ohhh that makes sense lol thank you. Psssst! *I'm combining rational expressions guys* x;